# Low-Altitude Wireless Networks: A Comprehensive Survey

### v2

@@ 6.2 Agentic AI-based LAWNs
Although lightweight large AI models substantially strengthen LAWNs by offering advanced perception and inference capabilities, the operational role of these models in current designs remains largely passive. In most architectures, large models are tied to predefined objectives, rely on offline pretraining, and function primarily as static function approximators that map inputs to outputs under fixed task specifications[Javaid2024]. The behavior is effectively constrained by the training data distribution and objective design, which limits the ability to autonomously revise goals, reinterpret evolving mission contexts, or sustain long-horizon coordination with other agents. Such rigidity is particularly problematic in LAWNs, where UAVs operate in nonstationary environments, face time-varying and sometimes conflicting mission objectives, and must continuously interact with other aircraft and ground infrastructure.
@@
Agentic AI offers a promising paradigm to overcome these limitations by elevating LAWNs from passive optimization engines to active, goal-directed autonomous systems. By embedding reasoning, memory, tool utilization, and self-evolution into network entities, agentic AI enables UAVs and ground infrastructure to adapt trajectories, allocate resources, and collaborate in swarms in a mission-aligned manner [sapkota2025uavs, gao2025agentic]. Through continual learning and shared intent representations, agentic LAWN systems are capable of mitigating performance drift induced by mobility and partial observability, thereby sustaining resilient closed-loop operation in complex low-altitude scenarios [acharya2025agentic]. In this context, large models and learning modules are no longer treated as static components; instead, they form the core of deliberative agents that can interpret goals, plan actions, and revise strategies online as the environment and mission requirements evolve.
@@
It is worth mentioning that, within the agentic paradigm, task integration emerges as the central organizing principle for AI-based LAWNs. Rather than treating sensing, communication, control, and cognition as loosely coupled functions, the network is structured around mission-level tasks that are decomposed into several interdependent subtasks, which are allocated to heterogeneous agents and continuously re-evaluated as environments change. In particular, a task-oriented cognitive layer can be adopted to interpret high-level intent and fuse multi-modal observations, while lower layers implement coordinated functions, e.g., ISAC, wireless networked control, and edge intelligence. By doing so, UAVs and ground nodes can reason jointly about trade-offs in communication reliability, control precision, and sensing accuracy to adapt behavior based on both local context and global performance metrics [feng2025multi, zhou2024comprehensive]. In this way, the mission goal is translated into consistent cross-layer decisions, followed by gradually reshaping task priorities and resource allocations based on real-time feedback from execution.
@@
Building on this integrated view, some recent research has begun to explore agentic optimization solutions that couple mathematical coordination frameworks with autonomous decision-making. Optimization-based approaches, originally formulated in a centralized manner, are extended to distributed and agent-based schemes in which each UAV acts as a decision-making agent handling local objectives and constraints while obeying global mission requirements. For example, distributed optimization has been employed to jointly design multi-role UAV functionality switching and trajectories for secure task offloading in UAV-assisted mobile edge computing, where local agents iteratively coordinate to satisfy communication and computation constraints under dynamic conditions [zhong2024distributed]. In parallel, LAWN-wide closed-loop designs have been formulated that integrate sensing, communication, and control into unified optimization problems, demonstrating that jointly tuned resource allocation and controller design can significantly enhance positioning accuracy and service reliability compared with decoupled baselines [wang2024semantic, zhou2024comprehensive]. Complementing optimization-driven coordination, agentic AI solutions build learning-based autonomy on top of this substrate. At the single-agent level, reinforcement learning (RL) has been widely adopted to endow individual UAVs with adaptive decision-making capabilities in complex, uncertain environments. Typical implementations rely on actor–critic architectures, deep Q-networks, or proximal policy optimization to learn control and path-planning policies directly from interaction with the environment [kaelbling1996reinforcement, konda1999actor, schulman2017proximal]. In soft actor–critic–based designs, for instance, a UAV agent observes its flight state and local environment and outputs continuous control actions that track reference trajectories while avoiding obstacles and following energy and communication constraints [cheng2020autonomous]. The entropy-regularized objective promotes a balance between exploitation of known good behaviors and exploration of new strategies, improving robustness to modeling errors and unforeseen disturbances. From the LAWN perspective, such a controller functions as a local decision engine that interprets mission-level commands and translates them into fine actuation in real time.
@@
When multiple UAVs and ground nodes need to coordinate, multi-agent reinforcement learning (MARL) extends this paradigm to networked settings, where centralized training and decentralized execution are applied[kopic2024collaborative]. Specifically, each UAV is modeled as an agent with partial local observations and a policy that must adapt under the nonstationary dynamics induced by other learning agents. To name a few, the authors in [cui2019multi] investigated MARL-based resource allocation schemes that enable UAVs to share spectral and power resources while jointly maximizing network throughput and fairness. Furthermore, the work of [liang2023multi] developed a multi-UAV collaborative search and attack framework that formulated exploration, target engagement, and collision avoidance in unknown environments as a multi-agent decision problem. In fact, such learned coordination strategies are intractable to engineer manually, highlighting the value of agentic learning in realizing cooperative, mission-aware behaviors in LAWNs, even though issues such as nonstationarity, credit assignment, and training stability remain challenging.
@@
A further step toward cognitively richer agentic AI integrates LLMs as reasoning engines for UAV agents. For example, LLM-based UAV path-planning frameworks have been proposed for industrial scenarios, where an LLM is employed to interpret mission descriptions, query external tools or simulators, and generate high-level plans that are subsequently translated into executable trajectories for individual UAVs [wang2024survey]. In addition, the studies in [Javaid2024, Li2025] provided a systematic overview of how LLMs can support intent understanding, dialogue-based supervision, and explainable decision-making in LAWN systems. Operational prototypes such as Net-GPT further demonstrated LLM-empowered intermediaries that maintain contextual information, enforce safety policies, and coordinate interactions among UAVs, human operators, and external services in real time [Piggott2023]. Collectively, these developments point to a transition toward cognitive agentic LAWNs, where LLMs handle high-level reasoning, task decomposition, and tool orchestration, while specialized learning or optimization modules are responsible for low-level control and communication under explicit safety and regulatory constraints.
@@ 7.1 LAA Structuring
Figure 7: The supported various LAWN services associated with different U-space stages based on [juntong2025low].
LAA is a fundamental element for the successful implementation of LAWNs, where the proper organization and structuring of airspace become crucial to ensure both operational efficiency and safety[jia2025hierarchical, 11159633]. Nevertheless, the corresponding configuration and structuring of LAA remain in the early stage. The performance and management efficiency of LAA will directly affect the operation of urban air mobility (UAM) systems, such that the LAA structure must not only accommodate uncrewed drones but also facilitate the integration with manned aircraft. According to the ICAO framework [FAA2022], the entire airspace is divided into categories A, B, C, D, E, and G, with the G category being non-controlled airspace, while the others are controlled airspace. LAA primarily corresponds to categories G and E, but it may partially overlap with other categories in emergencies. The ultimate goal of LAA planning is to allow for the joint operation of low-altitude aircraft and traditional aviation. Against this background, the SESAR program [SESAR2017] defined a four-stage (U1-U4) development strategy for LAA, which is particularized as:
@@
Along with this evolution, the service sets associated with U1 through U4 are summarized in Fig. 7. In contrast to high-altitude airspace, LAA is significantly influenced by terrestrial environments, e.g., available flight altitudes, ground obstacles, aircraft noise, and weather conditions [Bauranov2021, Tang2022]. Generally, these factors can be categorized into safety, environmental, and social concerns. Among them, safety is the most paramount factor in low-altitude operations with meteorological conditions and ground obstacles being particularly critical. For instance, adverse weather such as strong winds/rain, and low temperatures can reduce aircraft endurance, impair battery life, and increase collision risks. Although weather forecasting and pre-planned trajectories can mitigate some unfavorable impacts, real-time predictions remain challenging. Moreover, environmental factors like noise and visual pollution also play a significant role in public acceptance of LAWNs, especially in urban areas. The work of [Schaffer2021] showed that drone noise has become a key concern, as the public is more sensitive to drone noise than to vehicle noise. In parallel, visual pollution is another significant issue when dense drones fly overhead, affecting public emotions. On top of this, social factors, particularly in terms of privacy and equality preservation, influence public acceptance as well. UAVs’ ability to capture personal data during operations raises privacy concerns, and low-altitude services are often perceived as a privilege for high-income groups, leading to resistance from lower-income groups [Sunil2015]. To address these challenges, existing research has structured LAA into five primary types based on various principles. i.e., pipeline airspace, corridor airspace, layered airspace, blocky airspace, and free airspace [juntong2025low]. Each type presents unique design challenges and advantages, and various studies have explored their feasibility, limitations, and applications.
@@
1) Pipeline Airspace: As shown in Fig. 8(a), pipeline airspace is the most commonly used form of structured airspace, typically characterized by the establishment of flight path structures or virtual three-dimensional zones within the LAA, akin to road networks on the ground. The pipeline does not necessarily adhere to a fixed geometric shape, and its width is generally designed to accommodate only a single aircraft at a time. Aircraft within the pipeline operate in a unidirectional flight mode and can adjust their altitude, heading, and other parameters at designated intersection points. Notably, pipeline airspace offers a low degree of operational freedom but ensures high safety levels. For instance, the Metropolis project in Europe, which serves as a representative model of pipeline airspace, has designed independent pipelines for both uncrewed and manned aircraft [metropolis2014]. The pipeline airspace consists of several key components: connector pipelines linking the ground to the lowest pipeline layer for takeoff and landing; horizontal pipelines connecting nodes within the same pipeline layer; and transition pipelines bridging different pipeline layers. When an aircraft passes through a marked node, it occupies that node for a specified period, with only one aircraft permitted to occupy the node at any given time [Sunil2015]. Each node maintains an interval list to track the expected occupancy duration of other aircraft. Airspace resources are allocated based on the sequence in which aircraft enter the node, and new aircraft are only allowed to pass through once the node’s time slot becomes available. It is important to note that the number of marked nodes in each pipeline layer increases as the altitude decreases for safety.
@@
2) Corridor Airspace: Compared to pipeline airspace, corridor airspace offers a broader coverage with fewer establishment requirements. As observed in Fig. 8(b), it is typically organized in a regular geometric shape, characterized by two key components in terms of a centerline and a width for meeting certain performance criteria. The corridor airspace allows multiple aircraft to fly in parallel following a bidirectional flight pattern. Moreover, corridors can be designed as single or multi-layered airspace to meet traffic flow requirements to increase LAWN capacity. Corridor airspace offers significant advantages such as fewer flight restrictions and greater resilience, where the horizontal and vertical separations among drones are ensured, allowing for maximum flight freedom under safe operational conditions [airbus2018]. To further unlock the potential of corridor airspace, new configurations are investigated to accommodate higher traffic demands. Embraer-X, a Brazilian aerospace company, has proposed dividing LAA into corridors, road segments, and boundary points. In particular, the corridors are dynamically determined by GPS waypoints to fulfill various air traffic flow requirements and improve system efficiency [embraerx2019]. The authors in [muna2021] introduced multi-layer air corridors, divided into top, middle, and bottom layers, in which the top and bottom layers are designated for drone flying purposes along north-south and east-west directions, respectively. While the middle layer serves as a transitional area for aircraft to temporarily wait and change directions, and thus effectively reducing the collision risk.
@@
3) Layered Airspace: Layered airspace refers to the division of LAA into several separate layers, each with the same or varying vertical intervals. Since layered airspace is primarily divided in the vertical direction, the operational rules of aircraft in each layer can be effectively controlled, and aircraft operating at different altitudes will not interfere with each other. It is noteworthy that this structured design enhances airspace safety and lowers the technical requirements at the expense of freedom of airspace utilization degradation. In response to the drone market requirements, Amazon proposed a layered LAA design for lightweight drones operating below 500 ft [amazon2015]. In particular, the framework consists of a low-speed traffic layer (below 200 ft) for non-transport missions or UAVs with limited onboard capabilities, a high-speed transport layer (200-400 ft) for faster UAVs capable of autonomous navigation and conflict avoidance in complex environments, and a no-fly zone (400-500 ft) restricted to emergency use, as depicted in Fig. 8(c).
@@
4) Blocky Airspace: Block-shaped airspace refers to the division of LAA into horizontal blocks to align with urban structures or specific functional requirements. These blocks differ in infrastructure technology and service provider configurations. Compared to corridor and layered airspace, block-shaped airspace offers greater flexibility, with fewer operational restrictions, allowing more efficient use of airspace. However, block-shaped airspace does not inherently provide aircraft separation, incurring significant challenges to effective airspace management and conflict avoidance. To achieve this, the SESAR CORUS project [corus2019] introduced X, Y, and Z block-shaped airspaces by jointly considering environmental/social factors and the offered service priority, as shown in Fig. 8(d). X airspace lies in the lowest service level with aircraft with low task risk, e.g., rural and sportful areas. By contrast, Y airspace is a higher level with strategic conflict avoidance services like those in urban areas. Z airspace is on the top, which supports larger traffic flows and mitigates extensive risk, e.g., the city center and near airports.
@@
5) Free Airspace: Different from the above structured LAA, free airspace is a non-structured framework that does not have fixed flight paths or divisions in either the horizontal or vertical direction. In the early 21st century, the National Aerospace Laboratory (NLR) of the Netherlands proposed free flight in the civil aviation domain [hoekstra2001]. The core objective is to enable aircraft to autonomously separate and avoid conflicts through the system, allowing enhanced freedom to select preferred flight paths. Based on this design, the Metropolis project [metropolis2014] proposed free LAA illustrated in Fig. 8(e). This mechanism only specifies the allowed flight altitude and takes responsibility for managing airspace capacity and setting airspace boundaries. The collision avoidance and aircraft separation are fully addressed by the drones themselves. Due to its unstructured nature, the capacity of free LAA is significantly enhanced. More importantly, the drones only face minimal flight restrictions, and thus significantly shorten their flight distances to the destinations. However, the successful application of free airspace requires a well-developed collision avoidance mechanism. As the missions become complicated, ensuring real-time collision avoidance in dense LAWN deployments remains an open issue. For ease of exposition, we summarize the main features of the aforementioned airspace types in TABLE 5, which provides a concise comparison by focusing on their structure, capacity, safety considerations, and a balanced overview of their advantages and disadvantages.

### v3

@@ 6.2 Agentic AI-based LAWNs
Although lightweight large AI models substantially strengthen LAWNs by offering advanced perception and inference capabilities, the operational role of these models in current designs remains largely passive. In most architectures, large models are tied to predefined objectives, rely on offline pretraining, and function primarily as static function approximators that map inputs to outputs under fixed task specifications[70]. The behavior is effectively constrained by the training data distribution and objective design, which limits the ability to autonomously revise goals, reinterpret evolving mission contexts, or sustain long-horizon coordination with other agents. Such rigidity is particularly problematic in LAWNs, where UAVs operate in nonstationary environments, face time-varying and sometimes conflicting mission objectives, and must continuously interact with other aircraft and ground infrastructure.
@@
Agentic AI offers a promising paradigm to overcome these limitations by elevating LAWNs from passive optimization engines to active, goal-directed autonomous systems. By embedding reasoning, memory, tool utilization, and self-evolution into network entities, agentic AI enables UAVs and ground infrastructure to adapt trajectories, allocate resources, and collaborate in swarms in a mission-aligned manner [136, 54]. Through continual learning and shared intent representations, agentic LAWN systems are capable of mitigating performance drift induced by mobility and partial observability, thereby sustaining resilient closed-loop operation in complex low-altitude scenarios [3]. In this context, large models and learning modules are no longer treated as static components; instead, they form the core of deliberative agents that can interpret goals, plan actions, and revise strategies online as the environment and mission requirements evolve.
@@
It is worth mentioning that, within the agentic paradigm, task integration emerges as the central organizing principle for AI-based LAWNs. Rather than treating sensing, communication, control, and cognition as loosely coupled functions, the network is structured around mission-level tasks that are decomposed into several interdependent subtasks, which are allocated to heterogeneous agents and continuously re-evaluated as environments change. In particular, a task-oriented cognitive layer can be adopted to interpret high-level intent and fuse multi-modal observations, while lower layers implement coordinated functions, e.g., ISAC, wireless networked control, and edge intelligence. By doing so, UAVs and ground nodes can reason jointly about trade-offs in communication reliability, control precision, and sensing accuracy to adapt behavior based on both local context and global performance metrics [51, 204]. In this way, the mission goal is translated into consistent cross-layer decisions, followed by gradually reshaping task priorities and resource allocations based on real-time feedback from execution.
@@
Building on this integrated view, some recent research has begun to explore agentic optimization solutions that couple mathematical coordination frameworks with autonomous decision-making. Optimization-based approaches, originally formulated in a centralized manner, are extended to distributed and agent-based schemes in which each UAV acts as a decision-making agent handling local objectives and constraints while obeying global mission requirements. For example, distributed optimization has been employed to jointly design multi-role UAV functionality switching and trajectories for secure task offloading in UAV-assisted mobile edge computing, where local agents iteratively coordinate to satisfy communication and computation constraints under dynamic conditions [203]. In parallel, LAWN-wide closed-loop designs have been formulated that integrate sensing, communication, and control into unified optimization problems, demonstrating that jointly tuned resource allocation and controller design can significantly enhance positioning accuracy and service reliability compared with decoupled baselines [161, 204]. Complementing optimization-driven coordination, agentic AI solutions build learning-based autonomy on top of this substrate. At the single-agent level, reinforcement learning (RL) has been widely adopted to endow individual UAVs with adaptive decision-making capabilities in complex, uncertain environments. Typical implementations rely on actor–critic architectures, deep Q-networks, or proximal policy optimization to learn control and path-planning policies directly from interaction with the environment [78, 84, 138]. In soft actor–critic–based designs, for instance, a UAV agent observes its flight state and local environment and outputs continuous control actions that track reference trajectories while avoiding obstacles and following energy and communication constraints [35]. The entropy-regularized objective promotes a balance between exploitation of known good behaviors and exploration of new strategies, improving robustness to modeling errors and unforeseen disturbances. From the LAWN perspective, such a controller functions as a local decision engine that interprets mission-level commands and translates them into fine actuation in real time.
@@
When multiple UAVs and ground nodes need to coordinate, multi-agent reinforcement learning (MARL) extends this paradigm to networked settings, where centralized training and decentralized execution are applied[85]. Specifically, each UAV is modeled as an agent with partial local observations and a policy that must adapt under the nonstationary dynamics induced by other learning agents. To name a few, the authors in [39] investigated MARL-based resource allocation schemes that enable UAVs to share spectral and power resources while jointly maximizing network throughput and fairness. Furthermore, the work of [96] developed a multi-UAV collaborative search and attack framework that formulated exploration, target engagement, and collision avoidance in unknown environments as a multi-agent decision problem. In fact, such learned coordination strategies are intractable to engineer manually, highlighting the value of agentic learning in realizing cooperative, mission-aware behaviors in LAWNs, even though issues such as nonstationarity, credit assignment, and training stability remain challenging.
@@
A further step toward cognitively richer agentic AI integrates LLMs as reasoning engines for UAV agents. For example, LLM-based UAV path-planning frameworks have been proposed for industrial scenarios, where an LLM is employed to interpret mission descriptions, query external tools or simulators, and generate high-level plans that are subsequently translated into executable trajectories for individual UAVs [154]. In addition, the studies in [70, 95] provided a systematic overview of how LLMs can support intent understanding, dialogue-based supervision, and explainable decision-making in LAWN systems. Operational prototypes such as Net-GPT further demonstrated LLM-empowered intermediaries that maintain contextual information, enforce safety policies, and coordinate interactions among UAVs, human operators, and external services in real time [127]. Collectively, these developments point to a transition toward cognitive agentic LAWNs, where LLMs handle high-level reasoning, task decomposition, and tool orchestration, while specialized learning or optimization modules are responsible for low-level control and communication under explicit safety and regulatory constraints.
@@ 7.1 LAA Structuring
Figure 7: The supported various LAWN services associated with different U-space stages based on [130].
LAA is a fundamental element for the successful implementation of LAWNs, where the proper organization and structuring of airspace become crucial to ensure both operational efficiency and safety[71, 72]. Nevertheless, the corresponding configuration and structuring of LAA remain in the early stage. The performance and management efficiency of LAA will directly affect the operation of urban air mobility (UAM) systems, such that the LAA structure must not only accommodate uncrewed drones but also facilitate the integration with manned aircraft. According to the ICAO framework [50], the entire airspace is divided into categories A, B, C, D, E, and G, with the G category being non-controlled airspace, while the others are controlled airspace. LAA primarily corresponds to categories G and E, but it may partially overlap with other categories in emergencies. The ultimate goal of LAA planning is to allow for the joint operation of low-altitude aircraft and traditional aviation. Against this background, the SESAR program [98] defined a four-stage (U1-U4) development strategy for LAA, which is particularized as:
@@
Along with this evolution, the service sets associated with U1 through U4 are summarized in Fig. 7. In contrast to high-altitude airspace, LAA is significantly influenced by terrestrial environments, e.g., available flight altitudes, ground obstacles, aircraft noise, and weather conditions [19, 149]. Generally, these factors can be categorized into safety, environmental, and social concerns. Among them, safety is the most paramount factor in low-altitude operations with meteorological conditions and ground obstacles being particularly critical. For instance, adverse weather such as strong winds/rain, and low temperatures can reduce aircraft endurance, impair battery life, and increase collision risks. Although weather forecasting and pre-planned trajectories can mitigate some unfavorable impacts, real-time predictions remain challenging. Moreover, environmental factors like noise and visual pollution also play a significant role in public acceptance of LAWNs, especially in urban areas. The work of [137] showed that drone noise has become a key concern, as the public is more sensitive to drone noise than to vehicle noise. In parallel, visual pollution is another significant issue when dense drones fly overhead, affecting public emotions. On top of this, social factors, particularly in terms of privacy and equality preservation, influence public acceptance as well. UAVs’ ability to capture personal data during operations raises privacy concerns, and low-altitude services are often perceived as a privilege for high-income groups, leading to resistance from lower-income groups [148]. To address these challenges, existing research has structured LAA into five primary types based on various principles. i.e., pipeline airspace, corridor airspace, layered airspace, blocky airspace, and free airspace [130]. Each type presents unique design challenges and advantages, and various studies have explored their feasibility, limitations, and applications.
@@
1) Pipeline Airspace: As shown in Fig. 8(a), pipeline airspace is the most commonly used form of structured airspace, typically characterized by the establishment of flight path structures or virtual three-dimensional zones within the LAA, akin to road networks on the ground. The pipeline does not necessarily adhere to a fixed geometric shape, and its width is generally designed to accommodate only a single aircraft at a time. Aircraft within the pipeline operate in a unidirectional flight mode and can adjust their altitude, heading, and other parameters at designated intersection points. Notably, pipeline airspace offers a low degree of operational freedom but ensures high safety levels. For instance, the Metropolis project in Europe, which serves as a representative model of pipeline airspace, has designed independent pipelines for both uncrewed and manned aircraft [119]. The pipeline airspace consists of several key components: connector pipelines linking the ground to the lowest pipeline layer for takeoff and landing; horizontal pipelines connecting nodes within the same pipeline layer; and transition pipelines bridging different pipeline layers. When an aircraft passes through a marked node, it occupies that node for a specified period, with only one aircraft permitted to occupy the node at any given time [148]. Each node maintains an interval list to track the expected occupancy duration of other aircraft. Airspace resources are allocated based on the sequence in which aircraft enter the node, and new aircraft are only allowed to pass through once the node’s time slot becomes available. It is important to note that the number of marked nodes in each pipeline layer increases as the altitude decreases for safety.
@@
2) Corridor Airspace: Compared to pipeline airspace, corridor airspace offers a broader coverage with fewer establishment requirements. As observed in Fig. 8(b), it is typically organized in a regular geometric shape, characterized by two key components in terms of a centerline and a width for meeting certain performance criteria. The corridor airspace allows multiple aircraft to fly in parallel following a bidirectional flight pattern. Moreover, corridors can be designed as single or multi-layered airspace to meet traffic flow requirements to increase LAWN capacity. Corridor airspace offers significant advantages such as fewer flight restrictions and greater resilience, where the horizontal and vertical separations among drones are ensured, allowing for maximum flight freedom under safe operational conditions [6]. To further unlock the potential of corridor airspace, new configurations are investigated to accommodate higher traffic demands. Embraer-X, a Brazilian aerospace company, has proposed dividing LAA into corridors, road segments, and boundary points. In particular, the corridors are dynamically determined by GPS waypoints to fulfill various air traffic flow requirements and improve system efficiency [47]. The authors in [121] introduced multi-layer air corridors, divided into top, middle, and bottom layers, in which the top and bottom layers are designated for drone flying purposes along north-south and east-west directions, respectively. While the middle layer serves as a transitional area for aircraft to temporarily wait and change directions, and thus effectively reducing the collision risk.
@@
3) Layered Airspace: Layered airspace refers to the division of LAA into several separate layers, each with the same or varying vertical intervals. Since layered airspace is primarily divided in the vertical direction, the operational rules of aircraft in each layer can be effectively controlled, and aircraft operating at different altitudes will not interfere with each other. It is noteworthy that this structured design enhances airspace safety and lowers the technical requirements at the expense of freedom of airspace utilization degradation. In response to the drone market requirements, Amazon proposed a layered LAA design for lightweight drones operating below 500 ft [12]. In particular, the framework consists of a low-speed traffic layer (below 200 ft) for non-transport missions or UAVs with limited onboard capabilities, a high-speed transport layer (200-400 ft) for faster UAVs capable of autonomous navigation and conflict avoidance in complex environments, and a no-fly zone (400-500 ft) restricted to emergency use, as depicted in Fig. 8(c).
@@
4) Blocky Airspace: Block-shaped airspace refers to the division of LAA into horizontal blocks to align with urban structures or specific functional requirements. These blocks differ in infrastructure technology and service provider configurations. Compared to corridor and layered airspace, block-shaped airspace offers greater flexibility, with fewer operational restrictions, allowing more efficient use of airspace. However, block-shaped airspace does not inherently provide aircraft separation, incurring significant challenges to effective airspace management and conflict avoidance. To achieve this, the SESAR CORUS project [37] introduced X, Y, and Z block-shaped airspaces by jointly considering environmental/social factors and the offered service priority, as shown in Fig. 8(d). X airspace lies in the lowest service level with aircraft with low task risk, e.g., rural and sportful areas. By contrast, Y airspace is a higher level with strategic conflict avoidance services like those in urban areas. Z airspace is on the top, which supports larger traffic flows and mitigates extensive risk, e.g., the city center and near airports.
@@
5) Free Airspace: Different from the above structured LAA, free airspace is a non-structured framework that does not have fixed flight paths or divisions in either the horizontal or vertical direction. In the early 21st century, the National Aerospace Laboratory (NLR) of the Netherlands proposed free flight in the civil aviation domain [61]. The core objective is to enable aircraft to autonomously separate and avoid conflicts through the system, allowing enhanced freedom to select preferred flight paths. Based on this design, the Metropolis project [119] proposed free LAA illustrated in Fig. 8(e). This mechanism only specifies the allowed flight altitude and takes responsibility for managing airspace capacity and setting airspace boundaries. The collision avoidance and aircraft separation are fully addressed by the drones themselves. Due to its unstructured nature, the capacity of free LAA is significantly enhanced. More importantly, the drones only face minimal flight restrictions, and thus significantly shorten their flight distances to the destinations. However, the successful application of free airspace requires a well-developed collision avoidance mechanism. As the missions become complicated, ensuring real-time collision avoidance in dense LAWN deployments remains an open issue. For ease of exposition, we summarize the main features of the aforementioned airspace types in TABLE 5, which provides a concise comparison by focusing on their structure, capacity, safety considerations, and a balanced overview of their advantages and disadvantages.